Mountain Dental Services is a specialized dental practice whose only service is filling cavities. Mountain has recorded the following for the past nine months:

Month Number of Cavities Filled Total Cost
January 300 $5,200
February 500 5,650
March 550 6,000
April 700 6,400
May 475 5,850
June 425 5,400
July 350 5,200
August 575 6,200
September 400 5,350

Required:

1. Use the high-low method to estimate total fixed cost and variable cost per cavity filled.

2. Using these estimates, calculate Mountain’s total cost for filling 450 cavities.

Presidio, Inc., produces one model of mountain bike. Partial information for the company follows:

Required:
1. Complete Presidio’s cost data table.
2. Calculate Presidio’s contribution margin ratio and its total contribution margin at each sales level indicated in the cost data table assuming the company sells each bike for $650.
3. Calculate net operating income (loss) at each of the sales levels assuming a sales price of $650.
Answer #1

1)Variable cost per cavity = Change in cost/Change in Cavity = (6400-5200)/(700-300) = 3 per cavity

Fixed cost = 6400-(700*3) = 4300

2) Total Cost = Fixed cost+Variable cost = 4300+(450*3) = 5650

Note: Please post each question individually as HOMEWORKLIB guidelines